..............The magnetic field strength of a current-carrying, circular loop along an x-axis running perpendicular to the loop and passing through its center is given by:

Bx = µ0IR2 2(x2 + R2)3/2

where R is the radius of the loop and I is the current. At what x position is Bx half the maximum found at the center of the loop?
Answer #1

By = Mo i R² 2 (R²+x2j3/2 at contre (x=o) B. Choose for Bree » 2R² = (p² + x²; 3/2 → (22)/3 = (2+x+y*%) » x²=(2213_1) R² x =